OpenStudy (anonymous):

An autographed baseball rolls off a 0.83m high desk and strikes the floor 0.64 m away from the desk. How fat was it rolling on the desk before it fell off? The acceleration of gravity es 9,81 m/s2 the answer in unites of m/s

OpenStudy (anonymous):

\[2x=2ut+at ^{2}\] \[u=o\] \[t =\pm \sqrt{2x \div a}\] that is the time it takes to fall if x=h so it moves at constant velocity horizontally so how fast does it have to be travelling to go 0.64m in that time?\[speed = distance \div time\]\[v _{horizontal} = x _{horizontal} \div (\sqrt{2h \div a})\]\[speed = 0.64 \div (\sqrt{2 \times 0.83 \div 9.81})\] \[speed \approx 1.56 ms ^{-1}\]